# Quantum chaos in $\mathcal{PT}$-symmetric quantum systems

### v1

@@ IV.2   OTOC
We now present our results from the calculation of the OTOC. More specifically, we present the OTOC calculations for various phases obtained from the CLSR analysis and across the transitions T1 and T2.
@@
$\mathcal{PT}$-symmetric integrable phase: The Hermitian case $\lambda=0$ has been studied previously  [42], where it was shown that the OTOC follows a power law for all time scales in the integrable phase. On the other hand, the OTOC features a transition from exponential growth at early times $t<t_{E}$ to a power law in the chaotic phase. In Fig. 4 (right), we show the OTOC $C(t)$ for the $\mathcal{PT}$-symmetric integrable phase with parameters $(\lambda,K)\in\{(0.001,0.1),(0.002,0.456),(0.03,1.135)\}$, and system-size $2^{14}$. We find that the OTOC exhibits a behavior similar to the Hermitian case, which agrees with the calculations of the mean CLSR.
@@
$\mathcal{PT}$-symmetric chaotic phase: Figure 4 (right) shows the OTOC for the $\mathcal{PT}$-symmetric chaotic phase for $\lambda=(10^{-5},10^{-4})$, and fixed $K=30$. As for the $\mathcal{PT}$-symmetric integrable phase, we find that the OTOC in the symmetric chaotic phase exhibits behavior similar to that of the Hermitian system in agreement with mean CLSR calculations.
@@
$\mathcal{PT}$-symmetry broken chaotic phase: This is the only $\mathcal{PT}$-symmetry broken phase present in the phase diagram. In Fig. 4 (left), we show the OTOC for $\lambda=(10^{-3},0.005,0.001)$, and fixed kicking strength $K=30$. In Fig. 4 (right), the OTOC for $(\lambda,K)\in\{(0.006,5.734),(0.01,15)\}$ is shown. It displays two separate regimes of exponential growth. In early times, the growth is dominated by the chaotic nature of the Hamiltonian. In contrast, at a late time, the growth is essential because of the complexity of eigenvalues due to $\mathcal{PT}$-symmetry breaking, as we explain below.
@@
Since, at late times, the evolution is dominated by the phase $e^{-iEt}$ from $E=E_{r}+I\alpha$. It follows that the late time norm $\mathcal{N}(t)$ obeys
@@
$\begin{split}&\hat{p}(0)\hat{p}(t)\hat{p}(t)\hat{p}(0)+\hat{p}(t)\hat{p}(0)\hat{p}(0)\hat{p}(t)\\ -&\hat{p}(t)\hat{p}(0)\hat{p}(t)\hat{p}(0)-\hat{p}(0)\hat{p}(t)\hat{p}(0)\hat{p}(t)\end{split}$ (12)
@@
In Figure 6, we show the Lyapunov exponent as well as the corresponding mean CLSR across the transition T2 for $\hbar_{\text{eff}}=(0.2,0.02)$. We find that for both the values of $\hbar_{\text{eff}}$, the Lyapunov exponent displays a peak. It is important to note that the peak does not appear at the transition T2 but instead exists inside the broken phase of the $\mathcal{PT}$ symmetry. A further investigation of the nature of the peak in the Lyapunov exponent profile (See Sec. V) is required.
We also not that in a calculation performed on the same system as ours [55], a different definition of the OTOC has been used, one involving both the operators $\theta$ and $p$. The OTOC defined this way displays polynomial growth in time in the $\mathcal{PT}$ symmetry broken regime and not an exponential ones as we observe.
Figure 6: The growth rate of $C(t)$ or equivalently, the Lyapunov exponent (LE) as well as the mean CLSR shows along the transition in T2 for $\hbar_{\text{eff}}=0.02,0.2$, calculated for system size $N=2^{12}$ and $K=30$. The Lyapunov exponent exhibits a peak for both the values of $\hbar_{\text{eff}}$ at a value of $\lambda$ which does not exactly coincide with the jump associated with CLSR.
V   Conclusion
@@
We obtain the Lyapunov exponent from a calculation of the OTOC. We show that the early time growth of the OTOC can be used to distinguish between the three phases we observe. In particular, once we define a normalized version of the OTOC to eliminate the effects of the growth in time due to the complex nature of eigenvalues, it can be employed to detect the transition from the $\mathcal{PT}$-symmetric chaotic phase to the $\mathcal{PT}$-symmetry broken chaotic phase. We also observe that the reality of the eigenvalues in the above two regimes does not display a sharp disappearance at the transition, and so the origin of the sharp jump in the CLSR, which indicates the transition between the two phases, needs to be investigated further.
Acknowledgements.

### v2

@@ 4.2 OTOC
We now present our results from the calculation of the OTOC. More specifically, we present the OTOC calculations for various phases obtained from the CLSR analysis and across the transitions $T_{1}$ and $T_{2}$.
@@
$\mathcal{PT}$-symmetric integrable phase: The Hermitian case $\lambda=0$ has been studied previously  [42], where it was shown that the OTOC follows a power law for all time scales in the integrable phase. On the other hand, the OTOC features a transition from exponential growth at early times $t<t_{E}$ to a power law in the chaotic phase. In Fig. 3 (right), we show the OTOC $C(t)$ for the $\mathcal{PT}$-symmetric integrable phase with parameters $(\lambda,K)\in\{(0.001,0.1),(0.002,0.456),\allowbreak(0.03,1.135)\}$, and system-size $2^{14}$. We find that the OTOC exhibits a behavior similar to the Hermitian case, which agrees with the calculations of the CLSR.
@@
$\mathcal{PT}$-symmetric chaotic phase: Figure 3 (right) shows the OTOC for the $\mathcal{PT}$-symmetric chaotic phase for $\lambda=(10^{-5},10^{-4})$, and fixed $K=30$. As for the $\mathcal{PT}$-symmetric integrable phase, we find that the OTOC in the symmetric chaotic phase exhibits behavior similar to that of the Hermitian system in agreement with CLSR calculations.
@@
$\mathcal{PT}$-symmetry broken chaotic phase: This is the only $\mathcal{PT}$-symmetry broken phase present in the phase diagram. In Fig. 3 (left), we show the OTOC for $\lambda=(10^{-3},0.005,0.001)$, and fixed kicking strength $K=30$. In Fig. 3 (right), the OTOC for $(\lambda,K)\in\{(0.006,5.734),(0.01,15)\}$ is shown. It displays two separate regimes of exponential growth. In early times, the growth is dominated by the chaotic nature of the Hamiltonian. In contrast, at late times, the growth is essentially because of the complexity of eigenvalues brought about by $\mathcal{PT}$-symmetry breaking, as we explain below.
@@
Since, at late times, the evolution is dominated by the phase $e^{-iEt}$ from $E=E_{r}+i\alpha$, it follows that the late time norm $\mathcal{N}(t)$ obeys
@@
$\begin{split}&\hat{p}(0)\hat{p}(t)\hat{p}(t)\hat{p}(0)+\hat{p}(t)\hat{p}(0)\hat{p}(0)\hat{p}(t)-\hat{p}(t)\hat{p}(0)\hat{p}(t)\hat{p}(0)-\hat{p}(0)\hat{p}(t)\hat{p}(0)\hat{p}(t)\end{split}$ (12)
@@
In Figure 6, we show the Lyapunov exponent as well as the corresponding CLSR across the transition $T_{2}$ for varying $\hbar_{\text{eff}}$. We find that for all values of $\hbar_{\text{eff}}$, the Lyapunov exponent displays a peak. It is important to note that the peak does not appear at the transition $T_{2}$ but instead exists inside the broken phase of the $\mathcal{PT}$ symmetry. A further investigation of the nature of the peak in the Lyapunov exponent profile (See Sec. 5) is required. We also note that in a calculation performed on the same system as ours [55], a different definition of the OTOC has been used, one involving both the operators $\theta$ and $p$. The OTOC defined this way displays polynomial growth in time in the $\mathcal{PT}$ symmetry broken regime and not an exponential one, as we observe.
Figure 5: The above figure shows a section of the CLSR contour map for different values of $\hbar_{\text{eff}}$. It can be observed that the transition T* moves closer to the integrable $\to$ chaotic transition seen in the classical model for which $K\sim 1$ as $\hbar_{\text{eff}}\to 0$. Furthermore, the PT-symmetric chaotic regime is observed to shrink in size with a decrease in $\hbar_{\text{eff}}$
4.3 Parameter dependence
We summarize here the behavior of the CLSR and OTOC as we take three particular limits of interest.
  • $N\to\infty$: This is the limit we take to cover the ring of allowed eigenvalues of $\hat{p}$ densely and approach the continuum limit. The CLSR and LE plots show little to no noticeable shift when increasing the value of $N$, as can be seen from Fig. 7. The phase transitions seen on the CLSR become sharper but still occur roughly at the same value of $\lambda$ with a slight drift.
  • $\lambda\to 0$: This limit signifies the Hermitian limit of the problem where only two regimes have been identified with the CLSR values 0.5 and 0.57 corresponding to the integrable and chaotic regimes, respectively. These correspond to the PT-symmetric integrable regime and PT-symmetric chaotic regime for the general $(\lambda\neq 0)$ case. It can be seen from Fig. 6 that the value of $K$ for the transition between the two phases in the Hermitian limit persists even for small values of $\lambda$ for finite values of $\hbar$.
  • $\hbar\to 0$: This limit maps the quantum problem to its classical counterpart where again, there exists a phase transition from the integrable phase to a chaotic phase. As can be seen in Fig. 6, as $\hbar\to 0$, the transition from integrable-PT symmetric to chaotic-PT symmetric is happening at $K\approx 1$, which is where the classical transition occurs [56]. This observation is also in agreement with the transition observed for the Hermitian quantum kicked rotor [57] as $\hbar\to 0$. Further, we observe the shrinking of the chaotic-PT symmetric regime as $\hbar\to 0$. From the Hermitian case, we know that the chaotic-PT symmetric regime will always exist on the $\lambda=0$ line. However, we postulate that in the limit as $\hbar\to 0$, the only phase present in the $\lambda\neq 0$ space will be the chaotic-PT unbroken phase, i.e, an infinitesimal value of $\lambda$ will be sufficient to break PT symmetry, leading to chaotic nature in this limit.
Figure 6: The above figure provides a comparison between the observed PT-symmetric chaotic $\to$ PT-symmetry broken chaotic regime transition probed by the CLSR and the accompanying LE values for the same set of parameters for $K$ and $\lambda$. These have been plotted for various values of $\hbar_{\text{eff}}$ that span three orders of magnitude. We observe that this transition observed by the CLSR is accompanied by a peak in the LE. Note that different values of $K$ have only been used to ensure the transition is properly captured, since, as previously stated, the $\mathcal{PT}$-symmetric chaotic regime shrinks as $\hbar_{\text{eff}}\to 0$.
Figure 7: The above figure plots the transition from the PT-symmetric integrable regime to the PT-symmetric chaotic regime for different system sizes that increase by a factor of $2$. $\hbar=0.2$ and $K=30$ for all the plots. It can be seen that the transition, as obtained from the CLSR gets sharper as the system size ($N$) increases. There is a slight shift in the value of $\lambda$ at which the transition occurs and this shift appears to decrease with successively increasing values of $N$. The peak in the LE also appears to remain at a nearly fixed value of $\lambda$ as the system size is varied.
5 Conclusions
We have performed a comprehensive numerical study of the non-Hermitian kicked rotor demonstrating, for the first time (to our knowledge), the presence of integrability and $\mathcal{PT}$ symmetry breaking transitions in a driven quantum model. The model we have studied exhibits three phases, i) A $\mathcal{PT}$ symmetric integrable phase, ii) A $\mathcal{PT}$ symmetric chaotic phase and iii) A $\mathcal{PT}$ symmetry broken chaotic phase, thus demonstrating that $\mathcal{PT}$ symmetry breaking is a sufficient condition for the onset of chaos. We have characterized these phases and the transitions between them by calculating the complex level spacing ratio (CLSR) and the out of time ordered correlator (OTOC).
@@ 5 Conclusions
We obtain the Lyapunov exponent from a calculation of the OTOC. We show that the early time growth of the OTOC can be used to distinguish between the three phases we observe. In particular, once we define a normalized version of the OTOC to eliminate the effects of the growth in time due to the complex nature of eigenvalues, it can be employed to detect the transition from the $\mathcal{PT}$-symmetric chaotic phase to the $\mathcal{PT}$-symmetry broken chaotic phase. We also observe that the reality of the eigenvalues in the above two regimes does not display a sharp disappearance at the transition, so the origin of the sharp jump in the CLSR, which indicates the transition between the two phases, needs further investigation.
Acknowledgements
@@
Appendix A Numerics for the level spacing calculations
This section describes the computational schemes employed along with the necessary theoretical motivation.
A.1 Modifying the Hamiltonian
The Hamiltonian of a $\mathcal{PT}$ symmetric kicked rotor is conventionally chosen to have the following form for the potential $V(\theta)$.
$V(\theta)=K(\cos{\theta}+i\lambda\sin{\theta})$ ) ) (16)
However, in order to be able to simultaneously study the effect of varying the kicking strength $K$ and the non-Hermiticity parameter $\lambda$, the above form is not a good choice. For $\lambda\gg 1$, we see that $\lambda$ not only tunes the non-Hermiticty parameter but also controls the overall ‘magnitude’ of the $V(\theta)$ term and so controls the kicking strength as well. To avoid this, we ‘normalize’ the term inside and use the following definition for $V(\theta)$.
$V(\theta)=\frac{K}{\sqrt{1+\lambda^{2}}}(\cos\theta+i\lambda\sin\theta)$ (17)
A.2 Accounting for degeneracies due to symmetries
While calculating the energy level spacing distribution from the eigenvalue spectrum of a Hamiltonian, it is very important that one accounts for degeneracies arising from symmetries of the Hamiltonian. Hence, it is a common practice to carry out the level spacing calculations on a subset of the set of eigenvalues without repeating degenerate eigenvalues. The Hermitian kicked rotor Hamiltonian has parity $\mathcal{P}$ as one of its symmetries. To achieve such a subset, we block diagonalize the Hamiltonian using a basis that diagonalizes the group of symmetries of the Hamiltonian. Then, we perform the level spacing calculations over one of the boxes of the Hamiltonian.
In our case, defining the Floquet operator and retrieving the eigenvalues results in a large number of unidentified symmetries and near symmetries, causing a peak to exist near $\Delta E$ (level spacing) $=0$ in the distribution, as can be seen in figure 8. To improve the statistics and get rid of such near symmetries, we modify the kinetic energy term in the Hamiltonian. When written in the angular momentum basis, the kinetic energy is diagonal with its diagonal elements being $-\dfrac{\hbar^{2}l^{2}}{2m}$ where $l$ is the angular momentum eigenvalue and $m$ the moment of inertia. To break such symmetries, we add randomness to the masses of each diagonal term individually by redefining the diagonal elements as $-\dfrac{\hbar^{2}l^{2}}{2(m+\delta m)}$. For our calculations, $m=1$ and the $\delta m$’s are chosen randomly from a uniform distribution over the interval $[0,10^{-3}]$ for each diagonal element. This has the desired effect while still preserving the transition from the integrable to the chaotic phase. Figure 8 provides an example of this in the Hermitian case. The Hermitian version of the PTKR model has $\mathcal{P}$ and $\mathcal{PT}$ as symmetries of the system. The plot is of the Wigner-Dyson level spacing distribution.
Figure 8: The level spacing distribution for $k\simeq 100$. It can be seen that due to degeneracies arising from symmetries or near-symmetries, the distribution does not match the GOE form (left panel). The distribution matches the GOE formafter introducing a random small perturbation of $10^{-3}m$ (right panel).
A.3 Unfolding procedure
Whether using real or complex eigenvalues, the unfolding process is crucial to getting a faithful representation of the energy level spacing distribution. The energy level spacing has the dimensions of energy. Wigner’s analysis of the level spacing distributions involved calculations over a large ensemble of $2\times 2$ matrices with particular symmetries. Suppose that this is also true for the distribution obtained from an $N\times N$ random matrix with the same symmetries over its spectrum. In this case, it must be that the level spacing in different parts of the spectrum is not related or weakly related. Consequently, in order to compare level spacing across distinct sections of the spectrum, it becomes necessary to work with dimensionless quantities. These quantities can be obtained by dividing the level spacing by the mean level spacing in a small vicinity surrounding them. Such a procedure is known as unfolding.
For a real spectrum, the mean is simply obtained by averaging the level spacing about that eigenvalue for a fixed radius of points. For a complex spectrum, we define $\frac{1}{\sqrt{\rho_{k}}}$ as the mean level spacing around an eigenvalues $\xi$. Here $\rho_{k}$ is defined as follows.
$\rho_{k}=\dfrac{3n}{\pi\left(r_{k,n-1}^{2}+r_{k,n}^{2}+r_{k,n+1}^{2}\right)}$ (18)
Above, $n$ is predefined depending on our Hilbert space dimensions $N$. In our calculations, $n=10$ appears to work well for $N=2001$. Note that real or complex level spacing ratios do not need to undergo unfolding procedures since they already involve averaging over dimensionless quantities.
A.4 Averaging over intervals of $K$
In order to make the level spacing ratios and distributions robust to any kind of special high symmetry $K$ values and to obtain better statistics at a much lower computational cost, we average over intervals of $K$. For kicking strengths far away from the transition points, the behavior of the energy level spacing ratio and distribution does not change much at all. Further, since the folding of the eigenvalue spectrum depends greatly on the kicking strength, there can be resonances in the quantum mechanical model reminiscent of those in the classical model. This can cause the suppression of the chaotic nature of the system. Also, rather than working with a higher dimension Hilbert space, which requires $\mathcal{O}(N^{2})$ time, we can simply combine the eigenvalues of nearby kicking strengths, which only takes $\mathcal{O}(N)$ time. Note that because the magnitude of eigenvalues can differ for different kicking strengths, we normalize the spectra for separate kicking strengths so that it lies in the interval $[0,1]$ and then concatenate them. This technique has not been employed for the calculations that yield color plots. However, this is very effective if one wishes to calculate the CLSR or RLSR in a particular neighborhood away from the transitions.
Appendix B Calculation on random matrix ensembles and new CLSR
In our work, we show the agreement of the level spacing ratio obtained from our model and that from calculations over random matrices with the same symmetries as those of the model. Some of these values have been calculated previously as well [PhysRevB.106.134202]. We perform random matrix calculations for complex universality classes GinUE, GinOE, and AI [49]. Here, we highlight the method used to perform such calculations.
First, a complex matrix with $2N^{2}$ real entries is defined, giving complex random matrices of size $N\times N$. Every real entry is independently and identically chosen from a normal distribution of mean 0 and standard deviation 1. Any further symmetry constraints are added as follows: Let the symmetry group of these constraints be $\mathcal{A}$, then all the elements of $\mathcal{A}$ are made to act on copies of the same random matrix, and all of the resultant matrices are added. Next, the eigenvalues of the matrix thus obtained are numerically calculated, and the CLSR is found. This process is repeated enough times so as to ensure that the standard deviation of all the measures is sufficiently small compared to the mean.
The complex level spacing ratio can also be defined using angles, which is the other definition provided in [50]. This definition has not been adopted by us since it fails for a real eigenspectrum where $\theta$ becomes ill-defined. However, we perform random matrix calculations for $-\langle\cos{\theta}\rangle$ ) ⟩ as well as described below.
Table 2: The above table consists of data gathered on the complex level spacing ratio for some complex universality classes (GinUE, GinOE, and $\text{AI}^{\dagger}$); bracket contains must be read as (mean value, standard deviation).
Universality class $\langle r\rangle$ $-\langle\cos{\theta}\rangle$ ) ⟩
A/GinUE (0.73809, 0.00207) (0.23297, 0.0093)
$\text{AI}^{\dagger}$ (0.7231, 0.0027) (0.18777, 0.0077)
Appendix C $\mathcal{PT}$-symmetry broken phases
The quasienergy of the PTKR model is complex i.e. $\epsilon=\epsilon_{r}+i\epsilon_{i}$. When the maximum value of the imaginary part (denoted by $\alpha$) exceeds a certain threshold value $\alpha_{c}$, we assume the $\mathcal{PT}$-symmetry to be broken. The yellow shaded region is the $\mathcal{PT}$ symmetry broken region. In parts of the region near the transitions, we obtain a CLSR close to 0.739. This is the value of the CLSR one obtains for the GinOE universality class, to which the Hamiltonian belongs. This is due to the fact that the Hamiltonian is not bound by Hermiticity, and it commutes with an anti-unitary operator (namely $\mathcal{PT}$). In the table below, we show calculations of the CLSR on random matrices in the GinOE class. Additionally, we also perform calculations on matrices within GinOE that commute with $\mathcal{PT}$. These are in agreement with the value of the CLSR we get in our phase diagram.
Table 3: The above table consists of data gathered on the CLSR for the $\mathcal{PT}$ symmetry broken chaotic regime and its comparison with CLSR in the complex universality classes GinOE and its subset of matrices that commute with $\mathcal{PT}$, bracket contents must be read as (mean value, standard deviation).
Universality class CLSR
Phase diagram (0.735)
GinOE (0.73809, 0.0048)
$\mathcal{PT}$-symmetric matrix (0.73942, 0.0054)